the sides of the square are extended to form a rectangle. one side of the square is extended by 5 cm and the other side by 2 cm. the area of the rectangle is 130 cm². Find the length at the side of the square.

Answer by rothauserc(4718)

let x be the length of a side of the square, then
130 = (x+5) * (x+2)
130 = x^2 +7x +10
x^2 +7x -120 = 0
factor the polynomial
(x+15)*(x-8) = 0
x = -15 or x = 8
we want the positive value for x
therefore, the length of a side of the square is 8 cm